# FLARE v2: A Recursive Framework for Program Comprehension Across Languages and Levels of Abstraction

### v1

@@ Abstract
Building on the classroom framework reported in heath2025flare, this paper proposes FLARE v2 as a recursive, semiotically informed account of how program meaning is constructed. It reinterprets the descriptive tiers of FLARE v1 as instances of a single generative operation: identify elements (characterised by the four properties Receives, Sends, Effects, Shares); analyse their bindings along two dimensions (Causal-Temporal and Communicative); and recognise the new element that emerges. The Causal-Temporal dimension encompasses three subtypes - Sequential, Branch, and Event - that together account for control flow in both procedural and event-driven environments. A Compositional Ladder provides a visual parallel between literacy progressions and programming structures, illustrating how recursive composition operates from blocks and statements through segments, systems, and services. The framework aims to address conceptual and cognitive-load limitations reported in FLARE v1 and is situated within semiotic and program-comprehension theory. FLARE v2 is presented as a conceptual lens with potential implications for pedagogy and curriculum design; implementation and empirical evaluation are left for future work.
@@ 1 Introduction
FLARE v1 (heath2025flare) responded to this need by adapting the Block Model for primary classrooms as four tiers for talking about block-based programs: Blocks, Segments, Relationships, and Macro. The CEP ’25 study showed that these tiers helped teachers structure questioning and classroom dialogue, but also exposed limitations. Teachers reported that the Relationships and Macro tiers were difficult to apply in simple programs and, at times, confusing; classroom analysis suggested that the Relationships tier in particular combined several distinct mechanisms, including temporal ordering, causal triggering, and shared state (heath2025flare). Introducing all four tiers at once also placed demands on teachers’ subject knowledge and working memory. This paper presents FLARE v2 as a recursive reformulation in which elements bind to form new elements at successive scales. The next section summarises the constraints that emerged in the CEP ’25 work before introducing the generative core of FLARE v2.
Conceptually, FLARE v2 can be read as an integrative extension of Pennington’s dual abstraction, Schulte’s Block Model, and Sorva’s account of notional machines in event-driven environments (pennington1987comprehension; schulte2008block; sorva2013notional). It aims to retain their insights while providing a compositional account of how program meanings are constructed across scales.
@@ 2.3 Cognitive Load and Simultaneous Introduction
These limitations prompted a reconceptualisation. Where FLARE v1 was primarily a pedagogical scaffold - a set of tiers to organise classroom talk - FLARE v2 aims to be a generative theoretical account of how program meaning is constructed at any scale.
@@ 3 Formalising the Pattern: Code as Semiotic System
The limitations identified in FLARE v1 indicate the need for a descriptive lens that applies uniformly across scales, separates mechanisms rather than bundling them, and supports progressive introduction. Teachers’ natural reasoning about code focused on sequencing, triggering, conditions, and shared state, yet FLARE v1 did not articulate these distinctions.
FLARE v2 is designed to address these challenges by grounding the framework in a light semiotic footing. Classical structural linguistics emphasises that meaning in language arises from relations within a system and from the composition of bounded units into larger structures (saussure1916course; hjelmslev1961prolegomena). FLARE v2 adopts only a minimal commitment from this tradition: that bounded units carry meaning, and that binding these units produces new ones. This perspective supports the recursive element-binding cycle developed in the subsequent sections.
Figure 1 illustrates how literacy and programming share a common compositional structure. On the left, familiar literacy units (words, sentences, paragraphs, texts) combine through grammatical bindings. On the right, programming constructs (literals/expressions, commands, segments, systems) follow the same recursive pattern. This parallel provides teachers with an accessible metaphor: just as sentences combine to form paragraphs that express complete ideas, commands combine to form segments that accomplish discrete computational tasks.
Figure 1: The Compositional Ladder: parallel compositional progressions in literacy and programming. Both domains build meaning through recursive binding of smaller units into larger ones.
@@ 3.1 Design origins
The redesign of FLARE began from practice rather than theory. While using FLARE v1 with non-specialist teachers in block-based environments that sit on a JavaScript event loop, certain programs repeatedly exposed gaps in the framework. Simple scripts that mix non-blocking timers with sequential commands often behave in ways that pupils and teachers find surprising. For example, a program that schedules "morning" to be printed after one second and then immediately prints "Good" will output Good morning, even though many learners predict the opposite. FLARE v1 described structure at the level of Blocks, Segments, Relations and Macro, but it did not give teachers a precise way to talk about this kind of timing behaviour, about non-blocking constructs, or about the event loop that underpins them. The Relations tier in particular blurred temporal order, causal triggering and shared state, and classroom transcripts showed the same mechanisms being invoked in both the Segments and Relations tiers.
@@
However, a further distinction became apparent when considering event-driven code. When a child writes an event handler such as button.onClick(...), the handler is registered, but invocation is controlled by the runtime’s event loop - an invisible while(true) that checks for events and dispatches handlers. Mechanically, this is just another branch inside a loop. But the child did not write that loop; they cannot see it. From the learner’s perspective, the handler ”just fires” when clicked. This suggested that event-driven control, while mechanically equivalent to branching, deserves separate treatment because the branching mechanism is hidden from the programmer.
A subsequent exploratory classroom probe (n=26, ages 9-10) revealed an interesting pattern. When shown event-driven Scratch-style code and asked what would happen if a second event fired mid-execution, 77% of pupils selected concurrent execution on the probe items, without prompts or worked examples in the session itself. These children had most recently encountered an event-handling metaphor two years earlier in Year 3, when we talked about indented code as ”levels” and they proposed an event-handling metaphor, noting that multiple handlers could work simultaneously. They had received no further teaching on concurrency in the intervening years. This suggests that such early metaphors may form lasting execution models, and that children in event-driven, block-based curricula can reason concurrently about event-driven code, even though the classical comprehension studies on which this work builds, such as Pennington’s FORTRAN studies, focused on sequential control-flow abstractions (pennington1987comprehension). Full methodology and results are reported in heath2025concurrency.
Pennington’s work on procedural FORTRAN documented how programmers in single-threaded languages reason about control-flow and data-flow within a sequential execution model (pennington1987comprehension). The finding that children reason concurrently without explicit instruction suggests that this empirical regularity may not generalise straightforwardly to learners whose first programming experiences are event-driven.
@@ 4 The Generative Core of FLARE v2
The Compositional Ladder and framing developed in Sections 2 and 3 suggest a simple generative move that recurs across scales: elements bind to form new elements, which can themselves be treated as units in further bindings. Figure 2 visualises this recursive cycle, and FLARE v2 formalises this move in two parts:
@@
    a set of binding dimensions that explain how elements combine to yield new bounded units.
Together these form a recursive element-binding-element cycle that applies to blocks, segments, systems, and larger structures.
Figure 2: The generative core of FLARE v2: elements at scale $n$ bind to form new elements at scale $n+1$. The same descriptive properties and binding mechanisms apply recursively at every level.
@@ 4.1 Elements and their four properties
Depending on context, an element may be a block, statement, function, script, subsystem, or larger component. At every scale, the same four observable properties apply:
@@
  • Effects: internal state changes confined within the element, such as updates to local variables or private data structures.
  • Shares: any cross-boundary state or resource that the element reads or writes, such as global variables, configuration values, device state, shared hardware pins, or external services.
The distinction between Effects and Shares is crucial. Effects are private and unobservable by other elements. Any change that can be observed outside the element - because it modifies shared state or interacts with hardware - is a Share. Effects describe private state; Shares describe public or external state. Transient messages or one-off calls that pass values between elements are treated as Sends, not Shares. Shares are reserved for persistent cross-boundary state or resources that can be read or written by multiple elements over time.
An important consequence of this distinction is that Effects do not participate in bindings between elements. Because Effects are private and unobservable, they cannot create relationships with other elements. Only Receives, Sends, and Shares generate the inter-element connections that constitute Communicative bindings. Effects matter for understanding what happens inside an element, but they do not contribute to how elements combine.
@@
Because the same four properties apply at every level, teachers can move seamlessly between discussing a single block and a whole program. In classroom use, these properties become question stems: ”What does this element receive?”, ”What does it send?”, ”What does it change inside itself?”, and ”What does it share?” The vocabulary remains the same regardless of scale.
Figure 3 shows how FLARE v2’s recursive structure can map across different programming paradigms and levels of abstraction. The framework applies equally to visual block-based languages, text-based languages, and even lower-level representations like assembly code. I argue that this apparent universality emerges because the element-binding pattern is a fundamental way to describe how computational meaning is constructed, regardless of surface syntax, at least in the imperative and event-driven contexts typical of school curricula.
Figure 3: FLARE v2 structural layers showing how elements exist at multiple scales, from sub-blocks through to ecosystems. The same recursive grammar applies at each level.
@@ 4.2 Binding dimensions: how elements combine
Section 2 identified that FLARE v1 conflated several distinct mechanisms under a single Relations tier. The design work described in Section 3 revealed that these mechanisms separate cleanly into two orthogonal dimensions, closely aligned with Pennington’s distinction between control-flow and data-flow abstractions (pennington1987comprehension).
FLARE v2 retains Schulte’s Block Model insight that comprehension involves structure, flow, and function (schulte2008block), but refactors these concerns into element properties (Receives, Sends, Effects, Shares) and binding dimensions (Causal-Temporal and Communicative). Two additions distinguish the present framework: recursion, which the Block Model does not formalise; and explicit treatment of Event binding, which accommodates event-driven paradigms that the Block Model’s imperative orientation does not directly address.
The dimensions are orthogonal in the sense that for any pair of elements, two independent questions can be asked: how does control move between them (Causal-Temporal), and how does information move (Communicative). The answers to these questions can vary independently.
@@ 4.2.1 Causal-Temporal binding: how control moves
The Causal-Temporal dimension describes how execution flows between elements. The assembly-level analysis in Section 3 revealed three distinct subtypes based on the underlying control mechanism:
@@
  • Event: The jump exists, but it is hidden inside the runtime’s event loop. When a child writes button.onClick(handler), they write only the handler; the dispatching loop that decides when to invoke it is invisible. Mechanically, Event reduces to Branch inside a loop - but from the learner’s perspective, the handler ”just fires” in response to external stimuli. The classification here is from the learner’s perspective: the hidden loop is precisely why Event warrants separate pedagogical treatment, even though it reduces mechanically to Branch.
The exploratory Year 5 probe reported in Section 3 found that 77% of pupils selected concurrent execution when a second event was triggered. These pupils had encountered an event-handling metaphor two years earlier in Year 3, and noted multiple handlers working at the same time, but they had received no further teaching on concurrency in the intervening years. This suggests that such early metaphors may support concurrent execution models that persist and transfer, rather than emerging only from later formal instruction.
@@ 4.2.2 Communicative binding: how data moves
The Communicative dimension describes how information flows between elements. It operates through three of the four element properties: Sends, Receives, and Shares. (Effects, being private, do not participate in inter-element binding.)
@@
  • Shares: Persistent state accessible to multiple elements - global variables, hardware state, configuration values.
Where the original FLARE v1 Relations tier bundled ”what triggers this” with ”what does this share”, the two-dimension scheme separates them cleanly. Causal-Temporal binding answers when and why control moves; Communicative binding answers what information moves and where it persists.
@@ 4.2.3 Why two dimensions, not three?
An earlier formulation of FLARE v2 proposed three binding dimensions: Causal-Temporal, Logical-Conditional, and Communicative-Referential. The Logical-Conditional dimension was intended to capture predicates, guards, and Boolean conditions. However, further analysis revealed that conditionals are not a separate kind of binding but rather a mechanism that implements branching control flow. The if block, the loop guard, and the Boolean expression are all constructs that produce Branch-subtype Causal-Temporal bindings.
Treating conditionals as a separate dimension conflated the mechanism (Boolean evaluation) with the effect (control transfer). The revised scheme places conditionals where they belong: as the means by which Branch bindings are created. This simplifies the framework without losing descriptive power. Questions about conditions (”Under what circumstances does this run?”) remain important but are now understood as questions about how a Branch binding is governed, not as a separate binding dimension.
@@
Event Jump hidden in runtime loop. What triggers this?
@@
These dimensions are not proposed as an exhaustive taxonomy; other patterns may matter in object-oriented, concurrent, or distributed systems. But these two dimensions with three Causal-Temporal subtypes have proved sufficient in my own practice for supporting reasoning and classroom talk in imperative and event-driven contexts typical of primary and lower-secondary classrooms.

### v2

# FLARE v2: A Recursive Framework for Program Comprehension Across Common Teaching Languages and Levels of Abstraction

@@ Abstract
Building on the classroom framework in Heath et al. (2025), this paper proposes FLARE v2 as a recursive, semiotically informed account of how program meaning can be described across abstraction scales in common teaching languages. It reframes FLARE v1’s tiers as one cycle: identify bounded elements (Receives, Sends, Effects, Shares), analyse bindings along two dimensions (Causal-Temporal and Communicative), and treat the bound set as a new element at the next scale. Causal-Temporal binding has three subtypes - Sequential, Branch, and Event - to distinguish user-authored control flow from event-driven control whose dispatch is hidden in the runtime. A Compositional Ladder visualises the same compositional move from blocks and statements through segments and systems.
FLARE v2 is scoped to imperative and event-driven environments typical of primary and lower-secondary curricula. Above the system layer, behaviour is increasingly shaped by interaction between code and operating context (scheduling, infrastructure, permissions, contracts, failures, platform policy). Here, the element-and-binding vocabulary remains a structural baseline, but continuity of explanation typically requires overlays that make environmental constraints explicit. Event binding and overlays serve a common pedagogical role - preserving coherent structural reasoning where key causal mechanisms are not fully visible in the authored artefact. OOP design reasoning, explicit concurrency models, distributed systems, and functional paradigms are treated as future extensions; implementation and evaluation are left for future work.
Keywords: program comprehension, semiotics, block-based programming, text-based programming, computing education, abstraction, recursive frameworks, formative assessment, classroom dialogue
@@ 1 Introduction
FLARE v1 (heath2025flare) responded to this need by adapting the Block Model for primary classrooms as four tiers for talking about block-based programs: Blocks, Segments, Relationships, and Macro. The CEP ’25 study showed that these tiers helped teachers structure questioning and classroom dialogue, but also exposed limitations. Teachers reported that the Relationships and Macro tiers were difficult to apply in simple programs and, at times, confusing; classroom analysis suggested that the Relationships tier in particular combined several distinct mechanisms, including temporal ordering, causal triggering, and shared state (heath2025flare). Introducing all four tiers at once also placed demands on teachers’ subject knowledge and working memory.
This paper presents FLARE v2 as a recursive reformulation in which elements bind to form new elements at successive scales. Section 2 states the intended contribution and boundaries of the claim. Section 3 summarises the constraints that emerged in the CEP ’25 work, before Sections 4 and 5 introduce the semiotic framing and the generative core of FLARE v2. A later positioning section relates FLARE v2 to established program comprehension models.
2 Scope and Intended Contribution
FLARE v2 is proposed as a conceptual and analytic framework for reasoning about program meaning across several levels of abstraction. Its contribution is theoretical: a compact descriptive scheme for explaining how bounded program elements can be treated as units, related through bindings, and re-described at increasing scales. The framework aims to provide a shared classroom-facing vocabulary that can support teacher-led analysis, classroom dialogue, and reflective interpretation of code behaviour - particularly in contexts where execution mechanisms are partially hidden from learners.
The framework is intentionally descriptive rather than explanatory in a psychological sense. It does not claim to model the cognitive processes by which learners construct understanding, nor does it assert that novices naturally reason in its terms. Where empirical observations are reported - such as the Year 5 probe in Section 4 - these motivate design choices and suggest patterns worth investigating, but do not validate FLARE v2 as a cognitive theory. Claims about pedagogical value are hypotheses for future study, not established effects.
The scope is bounded by programming paradigm and by where meaning can plausibly be treated as largely ”in the artefact” for classroom analysis. FLARE v2 has been developed for imperative and event-driven environments typical of primary and lower-secondary computing curricula, including block-based and hybrid languages. The element and binding vocabulary is intended to remain usable as scale increases from expressions and commands through segments and systems.
Beyond this, service- and ecosystem-level reasoning increasingly requires accounting for factors that sit outside the authored artefact, so continuity of explanation depends on explicitly pairing code-side structure with environmental constraints. In such settings, the behaviour of an otherwise stable code artefact may depend on non-local factors (runtime scheduling, infrastructure configuration, permissions, service contracts, versioning, latency, partial failure, and platform policy). This does not require abandoning a structural-compositional lens, but it does mean that a purely code-centred structural description is no longer sufficient on its own. In system-plus contexts, FLARE v2 should be treated as describing the code-side composition, and paired with additional overlays that represent operational and socio-technical constraints.
Accordingly, this paper does not claim demonstrated applicability to object-oriented design reasoning, explicit concurrency models, distributed systems, or functional paradigms. Where such paradigms are discussed, they are treated as potential extensions rather than covered cases. In this sense, FLARE v2 is best understood as a candidate analytic lens: its success should be judged by whether it proves generative for analysis, curriculum design, and subsequent empirical investigation.
@@ 3.3 Cognitive Load and Simultaneous Introduction
These limitations prompted a reconceptualisation. Where FLARE v1 was primarily a pedagogical scaffold - a set of tiers to organise classroom talk - FLARE v2 aims to be a generative theoretical account of how program meaning can be described at increasing scale.
@@ 4 Formalising the Pattern: Code as Semiotic System
The limitations identified in FLARE v1 indicate the need for a descriptive lens that separates mechanisms rather than bundling them, and supports progressive introduction. Teachers’ natural reasoning about code focused on sequencing, triggering, conditions, and shared state, yet FLARE v1 did not articulate these distinctions.
FLARE v2 is designed to address these challenges by grounding the framework in a light semiotic footing. Classical structural linguistics emphasises that meaning in language arises from relations within a system and from the composition of bounded units into larger structures (saussure1916course; hjelmslev1961prolegomena). FLARE v2 adopts only a minimal commitment from this tradition: that bounded units carry meaning, and that binding these units yields larger units that can be re-described. This perspective supports the recursive element-binding cycle developed in the subsequent sections.
Figure 1 illustrates how literacy and programming can be discussed using a common compositional metaphor. On the left, familiar literacy units (words, sentences, paragraphs, texts) combine through grammatical bindings. On the right, programming constructs (literals/expressions, commands, segments, systems) can be treated using an analogous compositional move. This parallel provides teachers with an accessible metaphor: just as sentences combine to form paragraphs that express complete ideas, commands combine to form segments that accomplish discrete computational tasks.
Figure 1: The Compositional Ladder: parallel compositional progressions in literacy and programming. Both domains build meaning through composition of bounded units into larger structures.
@@ 4.1 Design origins
The redesign of FLARE began from practice rather than theory. While using FLARE v1 with non-specialist teachers in block-based environments that sit on a JavaScript event loop, certain programs repeatedly exposed gaps in the framework. Simple scripts that mix non-blocking timers with sequential commands often behave in ways that pupils and teachers find surprising. For example, a program that schedules "morning" to be printed after one second and then immediately prints "Good" will output Good morning, even though many learners predict the opposite. FLARE v1 described structure at the level of Blocks, Segments, Relations and Macro, but it did not give teachers a precise way to talk about timing behaviour, non-blocking constructs, or the event loop that underpins them. The Relations tier in particular blurred temporal order, causal triggering and shared state.
@@
A further distinction became apparent when considering event-driven code. When a child writes an event handler such as button.onClick(...), the handler is registered, but invocation is controlled by the runtime’s event loop - an invisible while(true) that checks for events and dispatches handlers. Mechanically, in common event-loop runtimes this can be approximated as Branch inside a loop. Pedagogically, the dispatch mechanism is hidden from the learner, and so the handler is experienced as something that ”just fires” when clicked. This motivated treating Event as a distinct subtype within Causal-Temporal binding in FLARE v2.
At this point, a tension emerges between the visible code and the observed runtime behaviour. In sequential and branching code, the learner can usually explain why one element runs next by pointing to a structure they authored: the next line, an if, a loop, or a call. In event-driven code, that causal link is no longer fully legible in the authored artefact. The learner writes a handler and registers it, but the decision about when it runs is made by a runtime dispatch mechanism that is typically invisible to them.
In FLARE v1, this loss of visible causality contributed to confusion within the Relationships tier, where temporal order, triggering, and shared state were discussed without a clean separation of mechanisms. FLARE v2 addresses this discontinuity by formalising the Event subtype within Causal-Temporal binding. The Event subtype functions as a pedagogical bridge: it reifies hidden dispatch into a distinct, nameable structural relation. By treating the event connection as a binding, the framework allows learners to maintain a structural mode of reasoning (“this element is bound to that trigger”) even when the underlying mechanism of that binding lies in the runtime environment.
On this view, Event binding does not claim that event-driven behaviour is unconstrained or fundamentally different from branching control. Rather, it marks a boundary of visibility and provides continuity in the learner’s explanatory model until they are ready to examine how the runtime implements dispatch (for example, via an event loop, callback queue, or scheduler).
A subsequent exploratory classroom probe (n=26, ages 9-10) revealed an interesting pattern. When shown event-driven Scratch-style code and asked what would happen if a second event fired mid-execution, 77% of pupils selected concurrent execution on the probe items, without prompts or worked examples in the session itself. These children had most recently encountered an event-handling metaphor two years earlier in Year 3, when we talked about indented code as ”levels” and they proposed an event-handling metaphor, noting that multiple handlers could work simultaneously. They had received no further teaching on concurrency in the intervening years. This suggests that early metaphors may form lasting expectations about event-driven execution, and it motivates careful treatment of scheduling and concurrency as an overlay when moving beyond single-threaded classroom runtimes. Full methodology and results are reported in heath2025concurrency.
@@ 5 The Generative Core of FLARE v2
The Compositional Ladder and framing developed in Sections 3 and 4 suggest a simple generative move that recurs across scales: elements bind to form new elements, which can themselves be treated as units in further bindings. Figure 2 visualises this recursive cycle, and FLARE v2 formalises this move in two parts:
@@
    a set of binding dimensions that describe how elements combine to yield new bounded units.
Together these form a recursive element-binding-element cycle that can be applied from blocks and statements through segments and systems, and extended to higher layers with additional overlays where required.
Figure 2: The generative core of FLARE v2: elements at scale $n$ bind to form new elements at scale $n+1$. The same descriptive questions recur across levels; at higher layers additional overlays are often required.
@@ 5.1 Elements and their four properties
Depending on context, an element may be a block, statement, function, script, subsystem, or larger component. Across the scales considered in this paper, the same four observable properties are used:
@@
  • Effects: internal state changes treated as confined within the element boundary, such as updates to local variables or private data structures.
  • Shares: any cross-boundary state or resource that the element reads or writes, such as global variables, configuration values, device state, shared hardware pins, shared objects, or external services.
The distinction between Effects and Shares is crucial. Effects are treated as private for inter-element reasoning. Any change that can be observed outside the element - because it modifies shared state, escapes via shared references, or interacts with hardware or external resources - is analysed as a Share. Transient messages or one-off calls that pass values between elements are treated as Sends, not Shares. Shares are reserved for persistent cross-boundary state or resources that can be read or written by multiple elements over time.
An important consequence of this distinction is that Effects are not ordinarily treated as participating in bindings between elements. Because Effects are analysed as private, they do not create inter-element connections. Inter-element Communicative bindings arise through Receives, Sends, and Shares. Effects matter for understanding what happens inside an element, but they do not, by definition, constitute a cross-boundary connection.
@@
Because the same four properties can be used at each scale in the covered scope, teachers can move between discussing a single block and a whole program using stable questions: ”What does this element receive?”, ”What does it send?”, ”What does it change inside itself?”, and ”What does it share?” The vocabulary remains consistent even as the grain size changes.
Figure 3 shows how FLARE v2’s recursive structure can map across different programming paradigms and levels of abstraction. The claim here is not that all paradigms are covered by the present scope, but that the element and binding vocabulary is intended to remain usable as the scale changes, with overlays where required.
Figure 3: The same element-binding cycle can be applied at each level, but higher layers typically require language- and context-specific overlays (for example, concurrency models, object identity and aliasing, and service-contract concerns).
@@ 5.2 Binding dimensions: how elements combine
Section 3 identified that FLARE v1 conflated several distinct mechanisms under a single Relations tier. The design work described in Section 4 suggested that these mechanisms separate cleanly into two orthogonal dimensions, aligned with Pennington’s distinction between control-flow and data-flow abstractions (pennington1987comprehension).
FLARE v2 retains Schulte’s Block Model insight that comprehension involves structure, flow, and function (schulte2008block), but refactors these concerns into element properties (Receives, Sends, Effects, Shares) and binding dimensions (Causal-Temporal and Communicative). Two additions distinguish the present framework: recursion, which the Block Model does not formalise; and explicit treatment of Event binding, which accommodates event-driven paradigms where control transfer is governed by hidden runtime mechanisms.
The dimensions are orthogonal in the sense that for any pair of elements, two independent questions can be asked: how does control move between them (Causal-Temporal), and how does information move (Communicative).
@@ 5.2.1 Causal-Temporal binding: how control moves
The Causal-Temporal dimension describes how execution flows between elements. FLARE v2 uses three subtypes:
@@
  • Event: The control transfer exists, but it is governed by a runtime dispatcher (for example an event loop) that is typically hidden from the learner. The learner writes the handler and registration, but not the dispatch mechanism that determines when invocation occurs. In common event-loop runtimes, this is often implementable as Branch inside a loop, but from the learner’s perspective the dispatch mechanism is not part of the written program and so warrants separate descriptive treatment.
@@ 5.2.2 Communicative binding: how data moves
The Communicative dimension describes how information flows between elements. It operates through three of the four element properties: Sends, Receives, and Shares. (Effects, being treated as private, do not ordinarily participate in inter-element binding.)
@@
  • Shares: Persistent state accessible to multiple elements - global variables, hardware state, configuration values, or shared resources.
Where the original FLARE v1 Relations tier bundled ”what triggers this” with ”what does this share”, the two-dimension scheme separates them. Causal-Temporal binding answers when and why control moves; Communicative binding answers what information moves and where it persists.
@@ 5.2.3 Why two dimensions, not three?
An earlier formulation of FLARE v2 proposed three binding dimensions: Causal-Temporal, Logical-Conditional, and Communicative-Referential. The Logical-Conditional dimension was intended to capture predicates, guards, and Boolean conditions. However, further analysis suggested that conditionals are not a separate kind of binding but a mechanism that governs branching control flow. The if block, the loop guard, and the Boolean expression are constructs that govern Branch-subtype Causal-Temporal bindings.
Treating conditionals as a separate dimension conflates the mechanism (Boolean evaluation) with the effect (control transfer). The revised scheme places conditionals where they belong: as the means by which Branch bindings are governed. Questions about conditions remain important but are understood as questions about how a Branch binding is controlled, not as a separate binding dimension.
@@
Event Control transfer governed by hidden runtime dispatch. What triggers this?
@@
These dimensions are not proposed as an exhaustive taxonomy; other patterns may matter in paradigms outside the scope stated in Section 2. The claim here is limited: within the imperative and event-driven environments typical of primary and lower-secondary classrooms, these two dimensions and three Causal-Temporal subtypes have been sufficient to support analysis and classroom talk in the author’s practice.